OpenStudy (anonymous):

Need help with another intergral. Never really done one like this before. tan^6(3y)sec^2(3y)dy

OpenStudy (anonymous):

fo u know c++

OpenStudy (anonymous):

\[\int\limits_{?}^{?}\tan ^6\left( 3y \right)\sec ^2(3y)dy\]

OpenStudy (anonymous):

I think I need to do u=tan(3y)? Maybe not even close? ,

OpenStudy (anonymous):

or u=tanx? So du=sec^2x?

OpenStudy (anonymous):

Anyone have any clue?

OpenStudy (anonymous):

turing test ..do u knw c++?

OpenStudy (turingtest):

yes, that will work, but it's\[u=\tan(3y)\]\[du=3\sec^2(3y)\]remember to use the chain rule all the way through @shiv no, sorry

OpenStudy (turingtest):

sorry\[u=\tan(3y)\]\[du=3\sec^2(3y)dy\]can't forget the dy! very important

OpenStudy (anonymous):

so would it be \[1/3\int\limits_{?}^{?}\tan(u)^6dy?\]

OpenStudy (turingtest):

no,\[u=\tan(3y)\]so after the sub there should be no tangent left

OpenStudy (anonymous):

oh... its \[1/3\int\limits_{?}^{?}u^6du\]

OpenStudy (turingtest):

yep :D

OpenStudy (anonymous):

you are a life saver. Can I just ask you one more quick here

OpenStudy (turingtest):

sure but I'm headed for dinner so gotta be quick

OpenStudy (anonymous):

ill try. \[\int\limits_{-1}^{-1/2}t^-2 \sin^2(1+1/2)dt sorry its t \to the -2 wont show up though\]

OpenStudy (anonymous):

Any direction helps. Do i have to set something u ,, then eventually change the limits of integration? I never know when to do that and feel this could be one..

OpenStudy (turingtest):

\[\int t^{-2}\sin^2?dt\]sorry, what is the integrand exactly?

OpenStudy (anonymous):

the end is sin^2(1+1/t)

OpenStudy (anonymous):

that 2 was wrong. you got the rest right. evaluated from -1 to -.5

OpenStudy (turingtest):

\[\int t^{-2}\sin^2(1+\frac1t)dt\]yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

evaluated from -1 to -1/2 if that changes anything

OpenStudy (turingtest):

\[u=1+\frac1t\]you don't have to change the bounds if you turn it back to t at the end, but if you want to evaluate it in terms of u you need to do this:\[u_1=1+\frac1{-1}=0\]\[u_2=1+\frac1{-1/2}=-1\]so those are your new limits (if you want to evaluate in terms of u the differential is\[du=-t^{-2}dt\]so the integral is\[-\int_{0}^{-1}\sin udu=\int_{-1}^{0}\sin u du\]Ok gotta eat, good luck!

OpenStudy (anonymous):

Thanks a lot! I appreicate it